OpenStudy (anonymous):

how do i factor 16a^2 - 4b^2

OpenStudy (cwrw238):

note that each term is a perfect square this is the difference of 2 squares so use the identity x^2 - y^2 = (x + y)(x - y) here x = 4a and can u say what y = ?

OpenStudy (anonymous):

4(2a+b)(2a-b)?

OpenStudy (cwrw238):

16a^2 - 4b^2 yes - you can simplify first by taking out 4: = 4 (4a^2 + b^2) = 4(2a + b)(2a - b)

OpenStudy (anonymous):

thank you

OpenStudy (cwrw238):

yw